# The adjoint state method for parametric definable optimization without smoothness or uniqueness

### 27 Mar 2026 v1

@@ 4.3 Conservative adjoint formula
This is our main result. We consider a slightly more general version of $D_{f}$ in (4.5), allowing for a general conservative field for $F$, beyond $\partial^{c}F$. The motivation for the proposed form of $D_{f}$ follows directly from the preceeding discussion in Section 4.2.
@@
It is now assumed that $F$ is a general locally Lipschitz function, not necessary differentiable. Recall that Assumptions 4.2() hold, that $D_{F}$ is a conservative field for $F$ and that the value function $f\colon\mathbb{R}^{q}\to\mathbb{R}$ is equal to
@@
$\theta\mapsto\Big\{u\in\mathbb{R}^{q}:\exists\,(x,\lambda,\mu)\in\mathrm{Sol_{pd}}(\theta),$
$(0,u)\in\operatorname*{conv\!}D_{F}(x,\theta)+\left(\operatorname{Jac}_{x}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\mu,\operatorname{Jac}_{\theta}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{\theta}\!H(x,\theta)^{\!\mathsf{T}}\mu\right)\Big\}.$
@@
The Clarke subdifferential is known to be conservative for definable functions [15]. Therefore, we can take the Clarke subdifferential of $F$ as the conservative $D_{F}$ for $F$ in Theorem 4.6() and everywhere $D_{F}$ is used. Optionally, for any $\theta\in\mathbb{R}^{q}$, we can replace $D_{F(\cdot,\theta)}$ by the Clarke partial subdifferential $\partial^{c}_{x}F(\cdot,\theta)$, which is contained in the projection of the joint sudifferential [13, Proposition 2.3.16].
The proof follows similar lines as [34, Theorem 12]. Define the set-valued map $D_{\mathcal{L}}\colon\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{p}\times\mathbb{R}^{q}\rightrightarrows\mathbb{R}^{n}\times\mathbb{R}^{q}$ as
$D_{\mathcal{L}}(y,\theta)$ $=D_{\mathcal{L}}(x,\lambda,\mu,\theta)$
$\coloneq D_{F}(x,\theta)+\left(\operatorname{Jac}_{x}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{x}\!H(x,\theta)^{\!\mathsf{T}}\mu,\operatorname{Jac}_{\theta}\!G(x,\theta)^{\!\mathsf{T}}\lambda+\operatorname{Jac}_{\theta}\!H(x,\theta)^{\!\mathsf{T}}\mu\right),$
where we use the shorthand notation $y=(x,\lambda,\mu)$. Note that $D_{\mathcal{L}}$ is a partial conservative field for $\mathcal{L}$, which ignores the dependency in $\lambda,\mu$. This will be possible thanks to complementarity and the structure of $\mathcal{L}$.
As $D_{\mathcal{L}}$, $\mathrm{Sol_{p}}$ and $\mathrm{Sol_{pd}}$ can be described by first-order formulas, they are definable; see [14, Theorem 1.13]. By definition, the definable conservative field $D_{F}$ is graph closed and compact valued and locally bounded. By the parametric Fermat’s rule in [36, Example 10.12], $D_{\mathcal{L}}$ has nonempty values. It is also immediate that $D_{\mathcal{L}}$ is definable, graph closed, locally bounded, and compact valued by Proposition 4.4.
According to [34, Theorem 2 and Lemma 6], to finish the proof, it remains to show that $D_{f}$ has the chain rule property along continuously differentiable and definable curves. So let $t\mapsto\theta(t)$ be a continuously differentiable and definable curve in $\mathbb{R}^{q}$. The following definable selections are possible using [39, paragraph 4.5].
We fix for now $t\mapsto y(t)\coloneq(x(t),\lambda(t),\mu(t))$ a definable selection in $t\rightrightarrows\mathrm{Sol_{pd}}(\theta(t))$, with $t\mapsto x(t)$ a definable selection in $t\rightrightarrows\mathrm{Sol_{p}}(\theta(t))$ and $t\mapsto(\lambda(t),\mu(t))$ a definable selection in $t\rightrightarrows\operatorname*{Mult}(x(t),\theta(t))$. Let $t\mapsto T(t)\coloneq(v(t),u(t))$ be a definable selection in $t\rightrightarrows D_{\mathcal{L}}(y(t),\theta(t))$. There is a definable selection $t\mapsto(z(t),w(t))$ in $t\rightrightarrows D_{F}(x(t),\theta(t))$ such that for all $t$
$v(t)$ $=z(t)+\operatorname{Jac}_{x}\!G(x(t),\theta(t))^{\!\mathsf{T}}\lambda(t)+\operatorname{Jac}_{x}\!H(x(t),\theta(t))^{\!\mathsf{T}}\mu(t)$
$u(t)$ $=w(t)+\operatorname{Jac}_{\theta}\!G(x(t),\theta(t))^{\!\mathsf{T}}\lambda(t)+\operatorname{Jac}_{\theta}\!H(x(t),\theta(t))^{\!\mathsf{T}}\mu(t).$
@@
are differentiable everywhere except at finitely many points, say at $t_{1},t_{2},\ldots,t_{M}$. By Proposition 4.4(), $\mathrm{Sol_{pd}}$ is locally bounded. It follows that $t\mapsto y(t)=(x(t),\lambda(t),\mu(t))$ has left and right limits everywhere and can be extended to an absolute continuous curve on $t_{\ell-1},t_{\ell}]$, for any $\ell\in\{2,\dots,M\}$.
Using the equalities in (4.4), the continuity of $f$ given by Lemma A.3(), and the continuity of $F$, $G$ and $H$, we have
@@
$\qquad+\lambda(t_{\ell}^{-})^{\!\mathsf{T}}G(x(t_{\ell}^{-}),\theta(t_{\ell}))+\mu(t_{\ell}^{-})^{\!\mathsf{T}}H(x(t_{\ell}^{-}),\theta(t_{\ell}))$
$\qquad-\lambda(t_{\ell-1}^{+})^{\!\mathsf{T}}G(x(t_{\ell-1}^{+}),\theta(t_{\ell-1}))-\mu(t_{\ell-1}^{+})^{\!\mathsf{T}}H(x(t_{\ell-1}^{+}),\theta(t_{\ell-1}))$
$=\int_{t_{\ell-1}}^{t_{\ell}}\frac{d}{dt}F(x(t),\theta(t))\,dt+\int_{t_{\ell-1}}^{t_{\ell}}\frac{d}{dt}\left[\lambda(t)^{\!\mathsf{T}}G(x(t),\theta(t))\right]dt$
$\qquad+\int_{t_{\ell-1}}^{t_{\ell}}\frac{d}{dt}\left[\mu(t)^{\!\mathsf{T}}H(x(t),\theta(t))\right]dt.$
@@
$f(\theta(t_{\ell}))-f(\theta(t_{\ell-1}))=\int_{t_{\ell-1}}^{t_{\ell}}\big[\langle{\dot{\theta}(t),w(t)}\rangle+\langle{\dot{x}(t),z(t)}\rangle\big]dt$
$\qquad+\int_{t_{\ell-1}}^{t_{\ell}}\big[\langle{\dot{\lambda}(t),G(x(t),\theta(t))}\rangle+\langle{\dot{\mu}(t),H(x(t),\theta(t))}\rangle\big]dt$
$\qquad+\int_{t_{\ell-1}}^{t_{\ell}}\big[\langle{\lambda(t),\operatorname{Jac}_{\theta}\!G(x(t),\theta(t))\dot{\theta}(t)}\rangle+\langle{\mu(t),\operatorname{Jac}_{\theta}\!H(x(t),\theta(t))\dot{\theta}(t)}\rangle\big]dt$
$\qquad+\int_{t_{\ell-1}}^{t_{\ell}}\big[\langle{\lambda(t),\operatorname{Jac}_{x}\!G(x(t),\theta(t))\dot{x}(t)}\rangle+\langle{\mu(t),\operatorname{Jac}_{x}\!H(x(t),\theta(t))\dot{x}(t)}\rangle\big]dt.$ (4.6)
We have that, for any $\ell\in\{2,\dots,M\}$, on $(t_{\ell-1},t_{\ell})$,
$t\mapsto\langle{\dot{\mu}(t),H(x(t),\theta(t)))}\rangle\equiv t\mapsto\langle{\dot{\lambda}(t),G(x(t),\theta(t))}\rangle\equiv 0.$ (4.7)
This follows from the fact that, for any $t\in(t_{\ell-1},t_{\ell})$, all components of $G(x(t),\theta(t))$ are zero except those corresponding to inactive constraints $g_{i}$, with $i\in\{1,\ldots,m\}\setminus I(x(t),\theta(t))$. Moreover, the corresponding components $\dot{\lambda}_{i}(t)$ of $\dot{\lambda}(t)$ are also zero. Indeed, by continuity of $G$, for every $s$ in some open interval $\mathcal{I}$ containing $t$, $g_{i}(\theta(s),x(s))<0$ and $\lambda_{i}(s)=0$ for any $i$ such that $g_{i}(\theta(t),x(t))<0$. Hence, $\dot{\lambda}_{i}(t)=0$ for all such $i$.
@@
Since $D_{\mathcal{L}}$ is definable and compact valued, Lemma A.4() says that there is a sequence of definable selections that is dense in $D_{\mathcal{L}}(y(t),\theta(t))$, for all $t$. Each element of the sequence satisfies (4.8) for almost all $t$. Combining the closure in Lemma A.4(), with the fact that a countable union of Lebesgue negligible set is Lebesgue negligible, we obtain, for almost every $t$,
@@
Then, we have, for almost every $t$,
@@
Note that the $t\mapsto y(t)$ has been fixed until now. Thanks to the density result in Lemma A.5(), we know that there exists a sequence of definable selections $\Big(\underbrace{(x_{\ell},\lambda_{\ell},\mu_{\ell})}_{=:y_{\ell}}\Big)_{\ell\in\mathbb{N}}$ such that, for all $\ell\in\mathbb{N}$ and all $t$, $y_{\ell}(t)\in\mathrm{Sol_{pd}}(\theta(t))$ and such that, for all $t$,
$D_{f}(\theta(t))$ $=\{u\in\mathbb{R}^{q}:\exists\,y\in\mathrm{Sol_{pd}}(\theta(t)),(0,u)\in D_{\mathcal{L}}(y,\theta(t))\}$
$=\operatorname*{cl}\bigcup_{\ell\in\mathbb{N}}\{u\in\mathbb{R}^{q}:(0,u)\in D_{\mathcal{L}}(y_{\ell}(t),\theta(t))\}.$
For each $\ell\in\mathbb{N}$, $y_{\ell}$ satisfies (4.9) which applies to an arbitrary selection $y(t)=(x(t),\lambda(t),\mu(t))\in\mathrm{Sol_{pd}}(\theta(t))$. Using the same density argument, it follows, that, for almost every $t$,
@@
4.4 Failure of the method without the definability assumption
Let us illustrate the importance of the definability assumptions. For this we demonstrate that the local Lipschitz continuity, path differentiability and even differentiability of the objective function $F$, together with the continuous differentiability of the constraint functions $G$ and $H$, are not sufficient for $D_{f}$ in Theorem 4.6() to be a conservative field for $f$. The restriction to a subclass of functions with additional properties is required. As already stated, in this work, we elected for definable functions that arise in many relevant applications.
Before presenting the result, we describe a fractal-type set that will be central to the proof. Indeed, we define the set $C\subset\mathbb{R}^{2}$ as $C\coloneq\cap_{\ell\in\mathbb{N}}C_{\ell}$, where for each $\ell\in\mathbb{N}$, $C_{\ell}$ is the union of $4^{\ell}$ squares of side length $1/4^{\ell}$ each and $(C_{\ell})_{\ell\in\mathbb{N}}$ form a nested decreasing sequence for the inclusion. In addition, the set $C$ has the following properties:
@@ 4.4 Failure of the method without the definability assumption
    $C$ is nonempty and closed with empty interior, and is included in $0,1]\times[0,1]$; hence it is compact;
@@
    $C$ projected on the first axis gives $0,1]\times\{0\}$, and projected on the second axis it gives $\{0\}\times[0,1]$.
Such a set $C$ exists and can be constructed; see [34, Section 4.1] for all the details and all its properties. An illustration of the construction [34, Figure 1] is reproduced exactly in Figure 1().
Figure 1: The fractal construction in [34, Section 4.1]. They start with the closed unit square in black. It is divided into 16 squares, each with a side length equal to one fourth of the original square’s side length. Only four of them at specific positions are kept and the others dropped. This process is repeated recursively on each square ad infinitum. The additional red lines represent projection of these sets on rotated axes. Considering $C_{\ell}$, $\ell\in\mathbb{N}$, the set obtained after $\ell$ steps ($C_{0}$ is the original square). We have that $C_{\ell+1}\subset C_{\ell}$ for all $\ell$. Then, let $C=\cap_{\ell\in\mathbb{N}}C_{\ell}$, which is closed. The projection of $C$ on each axes are full segments. Furthermore, in the limit, both projections on rotated axes are Cantor sets of zero measure.
@@
    is differentiable and path differentiable on $\mathbb{R}^{2}$, continuously differentiable on $\mathbb{R}^{2}\setminus C$, and
@@
    is such that the set-valued map $D_{f}\colon\mathbb{R}\to\mathbb{R}$ given by
@@
By Claim A.7(), there exists a differentiable, locally Lipschitz and path differentiable function $F\colon\mathbb{R}^{2}\to\mathbb{R}$ such that $F(x,\theta)=0$ for every $(x,\theta)\in C$, and $F(x,\theta)>0$ for every $(x,\theta)\in\mathbb{R}^{2}\setminus C$. In addition, $F$ is continuously differentiable everywhere except on $C$ with
$\mathcal{B}_{c}(0,1)\subset\partial^{c}F(x,\theta)\quad\forall\,(x,\theta)\in C.$ (4.10)
Given the fact that $\operatorname*{arg\,min}_{(x,\theta)\in\mathbb{R}^{2}}F(x,\theta)=C$ is not definable, $F$ is not definable either.
@@
[…]
@@
Thanks to Definition $4\tagform@ii$(), for every $\theta\in[0,1]$, there exists $x\in[0,1]$ such that $(x,\theta)\in C$. It follows that the value function is zero on $[0,1]$, i.e,
@@
Let $\theta\in[0,1]$ and let $\bar{x}\in\mathrm{Sol_{p}}(\theta)$. As $\bar{x}\in[0,1]$ and $G$ is therefore not active at $(\bar{x},\theta)$, we have $(\bar{x},0)\in\mathrm{Sol^{c}_{pd}}(\theta)$. It follows that
$D_{f}(\theta)$ $=\big\{u\in\mathbb{R}:\exists\,(x,\lambda)\in\mathrm{Sol^{c}_{pd}}(\theta),(0,u)\in\partial^{c}F(x,\theta)+\left(\lambda\nabla_{\!x}G(x,\theta),0\right)\big\}$
@@
Consider the absolutely continuous curve $\gamma\colon[0,1]\to\mathbb{R}$ defined as $t\mapsto t$. Let’s now suppose, for the sake of contradiction, that $f$ has the chain rule property for $D_{f}$. Then, for almost all $t\in[0,1]$, and for any $u\in D_{f}(\gamma(t))$, in particular any $u\in[-1,1]$,
@@
That’s impossible. So $f$ does not have the chain rule property for $D_{f}$, i.e, $D_{f}$ is not conservative for $f$. ∎

### 22 Apr 2026 v2

@@ 4.3 Conservative adjoint formula
This is our main result. We consider a slightly more general version of $D_{f}$ in (4.7), allowing for a general conservative field for $F$, beyond $\partial^{c}F$. The motivation for the proposed form of $D_{f}$ follows directly from the preceding discussion in Section 4.2().
@@
Let Assumptions 4.4() hold. Recall that $F$ is a general locally Lipschitz function, not necessary differentiable, $D_{F}$ is a conservative field for $F$ and the value function $f\colon\mathbb{R}^{q}\to\mathbb{R}$ is the mapping
@@
$\theta\mapsto$ $\big\{u\in\mathbb{R}^{q}:\exists\,(x,\lambda)\in\mathrm{Sol_{pd}}(\theta),(0,u)\in\operatorname*{conv\!}D_{F}(x,\theta)+\operatorname{Jac}C(x,\theta)^{\!\mathsf{T}}\lambda\big\}.$
@@
The Clarke subdifferential is known to be conservative for definable functions [18]. Therefore, we can take the Clarke subdifferential of $F$ as the conservative $D_{F}$ for $F$ in Theorem 4.9() and everywhere $D_{F}$ is used. Optionally, for any $\theta\in\mathbb{R}^{q}$, we can replace $D_{F(\cdot,\theta)}$ by the Clarke partial subdifferential $\partial^{c}_{x}F(\cdot,\theta)$, which is contained in the projection of the joint subdifferential [15, Proposition 2.3.16].
The proof follows similar lines as [40, Theorem 12]. Define the set-valued map $D_{\mathcal{L}}\colon\mathbb{R}^{n}\times\mathbb{R}^{m}\times\mathbb{R}^{q}\rightrightarrows\mathbb{R}^{n}\times\mathbb{R}^{q}$ as
$D_{\mathcal{L}}(y,\theta)$ $=D_{\mathcal{L}}(x,\lambda,\theta)\coloneq D_{F}(x,\theta)+\left(\operatorname{Jac}_{x}\!C(x,\theta)^{\!\mathsf{T}}\lambda,\operatorname{Jac}_{\theta}\!C(x,\theta)^{\!\mathsf{T}}\lambda\right),$
where we use the shorthand notation $y=(x,\lambda)$. Note that $D_{\mathcal{L}}$ is a partial conservative field for $\mathcal{L}$, which ignores the variations in $\lambda\in\mathbb{R}^{m}$. This will be possible thanks to complementarity and the structure of $\mathcal{L}$.
As $D_{\mathcal{L}}$, $\mathrm{Sol_{p}}$ and $\mathrm{Sol_{pd}}$ can be described by first-order formulas involving only definable objects, they are definable; see [17, Theorem 1.13]. By definition, the definable conservative field $D_{F}$ is graph closed, compact valued and locally bounded. By the parametric Fermat’s rule in [42, Example 10.12], $D_{\mathcal{L}}$ has nonempty values. It is also immediate that $D_{\mathcal{L}}$ is definable, graph closed, locally bounded, and compact valued by Proposition 4.6().
According to [40, Theorem 2 and Lemma 6], to finish the proof, it remains to show that $D_{f}$ has the chain rule property along continuously differentiable and definable curves. So let $t\mapsto\theta(t)$ be a continuously differentiable and definable curve in $\mathbb{R}^{q}$. The following definable selections are possible using [46, paragraph 4.5].
We fix for now $t\mapsto y(t)\coloneq(x(t),\lambda(t))$ a definable selection in $t\rightrightarrows\mathrm{Sol_{pd}}(\theta(t))$, with $t\mapsto x(t)$ a definable selection in $t\rightrightarrows\mathrm{Sol_{p}}(\theta(t))$ and $t\mapsto\lambda(t)$ a definable selection in $t\rightrightarrows\operatorname*{Mult}(x(t),\theta(t))$. Let $t\mapsto T(t)\coloneq(v(t),u(t))$ be a definable selection in $t\rightrightarrows D_{\mathcal{L}}(y(t),\theta(t))$. There is a definable selection $t\mapsto(z(t),w(t))$ in $t\rightrightarrows D_{F}(x(t),\theta(t))$ such that for all $t$
$v(t)$ $=z(t)+\operatorname{Jac}_{x}\!C(x(t),\theta(t))^{\!\mathsf{T}}\lambda(t)$
$u(t)$ $=w(t)+\operatorname{Jac}_{\theta}\!C(x(t),\theta(t))^{\!\mathsf{T}}\lambda(t).$
@@
are differentiable everywhere except at finitely many points, say at $t_{1},t_{2},\ldots,t_{M}$. By Proposition 4.6(), $\mathrm{Sol_{pd}}$ is locally bounded. It follows from the monotonicity Lemma [17, Section 2.1]that $t\mapsto y(t)=(x(t),\lambda(t))$ has left and right limits everywhere and can be extended to an absolutely continuous curve on $t_{\ell-1},t_{\ell}]$, for any $\ell\in\{2,\dots,M\}$.
Using the equalities in (4.6), the continuity of $f$ given by Lemma A.3(), and the continuity of $F$ and $C$, we have
@@
$\qquad+\lambda(t_{\ell}^{-})^{\!\mathsf{T}}C(x(t_{\ell}^{-}),\theta(t_{\ell}))-\lambda(t_{\ell-1}^{+})^{\!\mathsf{T}}C(x(t_{\ell-1}^{+}),\theta(t_{\ell-1}))$
$=\int_{t_{\ell-1}}^{t_{\ell}}\frac{d}{dt}F(x(t),\theta(t))\,dt+\int_{t_{\ell-1}}^{t_{\ell}}\frac{d}{dt}\left[\lambda(t)^{\!\mathsf{T}}C(x(t),\theta(t))\right]dt.$
@@
$f(\theta(t_{\ell}))-f(\theta(t_{\ell-1}))=\int_{t_{\ell-1}}^{t_{\ell}}\big[\langle{\dot{\theta}(t),w(t)}\rangle+\langle{\dot{x}(t),z(t)}\rangle\big]dt+\int_{t_{\ell-1}}^{t_{\ell}}\big[\langle{\dot{\lambda}(t),C(x(t),\theta(t))}\rangle\big]dt$
$+\int_{t_{\ell-1}}^{t_{\ell}}\big[\langle{\lambda(t),\operatorname{Jac}_{\theta}\!C(x(t),\theta(t))\dot{\theta}(t)}\rangle+\langle{\lambda(t),\operatorname{Jac}_{x}\!C(x(t),\theta(t))\dot{x}(t)}\rangle\big]dt.$ (4.8)
We note that for all $t,C(x(t),\theta(t))\in K$ and $\lambda(t)\in N_{K}(C(x(t),\theta(t))$ thanks to Lemma 4.2(). Since both curves are continuously differentiable on $(t_{\ell-1},t_{\ell})$, Lemma 4.8() ensures that
$t\mapsto\langle{\dot{\lambda}(t),C(x(t),\theta(t))}\rangle\equiv 0\;\text{ on }(t_{\ell-1},t_{\ell}).$ (4.9)
@@
Since $D_{\mathcal{L}}$ is definable and compact valued, Lemma A.4() says that there is a sequence of definable selections, as $v$ and $u$, that is dense in $D_{\mathcal{L}}(y(t),\theta(t))$, for all $t$. Each element of the sequence satisfies (4.10) for almost all $t$. Combining the closure in Lemma A.4(), with the fact that a countable union of Lebesgue negligible set is Lebesgue negligible, we obtain, for almost every $t$,
@@
Then, using the fact that $(0,u)\in D_{\mathcal{L}}(y(t),\theta(t))$ for all $t$, we have, for almost every $t$,
@@
Note that the $t\mapsto y(t)$ has been fixed until now. Thanks to the density result in Lemma A.5(), we know that there exists a sequence of definable selections $\Big(\underbrace{(x_{\ell},\lambda_{\ell})}_{=:y_{\ell}}\Big)_{\ell\in\mathbb{N}}$ such that, for all $\ell\in\mathbb{N}$ and all $t$, $y_{\ell}(t)\in\mathrm{Sol_{pd}}(\theta(t))$ and such that, for all $t$,
$D_{f}(\theta(t))$ $=\big\{u\in\mathbb{R}^{q}:\exists\,y\in\mathrm{Sol_{pd}}(\theta(t)),(0,u)\in D_{\mathcal{L}}(y,\theta(t))\big\}$
$=\operatorname*{cl}\bigcup_{\ell\in\mathbb{N}}\big\{u\in\mathbb{R}^{q}:(0,u)\in D_{\mathcal{L}}(y_{\ell}(t),\theta(t))\big\}.$
For each $\ell\in\mathbb{N}$, $y_{\ell}$ satisfies (4.11) which applies to an arbitrary selection $y(t)=(x(t),\lambda(t))\in\mathrm{Sol_{pd}}(\theta(t))$. Using the same density argument, it follows, that, for almost every $t$,
@@
4.4 Gradient-like algorithms using the adjoint method
The main result of the previous section ensures that the adjoint method provides a selection in a conservative field for $f$. General conservative fields can be used as a first order oracle to optimize $f$ [9]. We describe below how the adjoint method yields an optimization method through a vanilla small-step gradient method for the problem:
$\min_{\theta\in\mathbb{R}^{q}}f(\theta).$
Proposition 4.11 (Small step method).
Under the assumptions of Theorem 4.9(), consider the recursion
$\theta_{k+1}=\theta_{k}-s\alpha_{k}\emph{ASM}(\theta_{k}),$
where ASM is a selection of $D_{f}$, $s>0$, $\alpha_{k}>0$ for all $k$ and $\alpha_{k}\to 0$ as $k\to\infty$. For almost all $\theta_{0}\in\mathbb{R}^{q}$ and all but finitely many $s>0$, if the sequence $(\theta_{k})_{k\in\mathbb{N}}$ is bounded, then any accumulation point $\bar{\theta}$ of $(\theta_{k})_{k\in\mathbb{N}}$ is Clarke critical, i.e., $0\in\partial^{c}f(\bar{\theta})$, where $\partial^{c}f(\bar{\theta})$ is the Clarke generalized gradient of $f$ at $\bar{\theta}$.
The proof is not provided since it is a direct consequence of [9, Theorem 9], [7, Theorem 7], see the proof of [40, Theorem 4].
Proposition 4.11() is related to the basic gradient method. The main device behind this result is continuous time Lyapunov decrease. Similar results hold for the heavy ball method [35], subgradient methods with Bregman distance [24] or adaptive methods such as Adam [48] under appropriate assumptions.
5 Failure of ASM without definability
Let us illustrate the importance of the definability assumptions in Problem (4.1). For this we demonstrate that the local Lipschitz continuity, path differentiability and even differentiability of the objective function $F$, together with the continuous differentiability of the constraint functions $G$ and $H$, are not sufficient for $D_{f}$ in Theorem 4.9() to be a conservative field for $f$. The restriction to a subclass of functions with additional properties is required. As already stated, in this work, we elected for definable functions that arise in many relevant applications.
Before presenting the result, we describe a fractal-type set that will be central to the proof. Indeed, we define the set $S\subset\mathbb{R}^{2}$ as $S\coloneq\cap_{\ell\in\mathbb{N}}S_{\ell}$, where for each $\ell\in\mathbb{N}$, $S_{\ell}$ is the union of $4^{\ell}$ squares of side length $1/4^{\ell}$ each and $(S_{\ell})_{\ell\in\mathbb{N}}$ form a nested decreasing sequence for the inclusion. In addition, the set $S$ has the following properties:
@@ 5 Failure of ASM without definability
    $S$ is nonempty and closed with empty interior, and is included in $0,1]\times[0,1]$; hence it is compact;
@@
    $S$ projected on the first axis gives $0,1]\times\{0\}$, and projected on the second axis it gives $\{0\}\times[0,1]$.
Such a set $S$ exists and can be constructed; see [40, Section 4.1] for all the details and all its properties. An illustration of the construction [40, Figure 1] is reproduced exactly in Figure 1().
Figure 1: The fractal construction in [40, Section 4.1]. They start with the closed unit square in black. It is divided into 16 squares, each with a side length equal to one fourth of the original square’s side length. Only four of them at specific positions are kept and the others dropped. This process is repeated recursively on each square ad infinitum. The additional red lines represent projection of these sets on rotated axes. Considering $S_{\ell}$, $\ell\in\mathbb{N}$, the set obtained after $\ell$ steps ($S_{0}$ is the original square). We have that $S_{\ell+1}\subset S_{\ell}$ for all $\ell$. Then, let $S=\cap_{\ell\in\mathbb{N}}S_{\ell}$, which is closed. The projection of $S$ on each axes are full segments. Furthermore, in the limit, both projections on rotated axes are Cantor sets of zero measure.
@@
    is differentiable and path differentiable on $\mathbb{R}^{2}$, continuously differentiable on $\mathbb{R}^{2}\setminus S$, and
@@
    is such that the set-valued map $D_{f}\colon\mathbb{R}\rightrightarrows\mathbb{R}$ given by
@@
By Claim A.7(), there exists a differentiable, locally Lipschitz and path differentiable function $F\colon\mathbb{R}^{2}\to\mathbb{R}$ such that $F(x,\theta)=0$ for every $(x,\theta)\in S$, and $F(x,\theta)>0$ for every $(x,\theta)\in\mathbb{R}^{2}\setminus S$. In addition, $F$ is continuously differentiable everywhere except on $S$ with
$\mathcal{B}_{c}(0,1)\subset\partial^{c}F(x,\theta)\quad\forall\,(x,\theta)\in S.$ (5.1)
Given the fact that $\operatorname*{arg\,min}_{(x,\theta)\in\mathbb{R}^{2}}F(x,\theta)=S$ is not definable, $F$ is not definable either.
@@
has the value $R(\theta)=[-2,2]$ for any $\theta\in\mathbb{R}$. Given $\theta\in\mathbb{R}$, the constraint is only active at $(-2,\theta)$ and $(2,\theta)$. Additionally, $\nabla_{\!x}G(-2,\theta)\neq 0$ and $\nabla_{\!x}G(2,\theta)\neq 0$; thus the MFCQ holds at any $(x,\theta)\in\operatorname*{gph}R$, i.e., Assumption $4\tagform@ii$() is satisfied.
@@
has nonempty values on $\mathbb{R}$, and $\mathrm{Sol_{p}}(\theta)\subset[-2,2]$ for any $\theta\in\mathbb{R}$, which means that $\mathrm{Sol_{p}}$ is uniformly bounded on $\mathbb{R}$; hence Assumption $4\tagform@iii$() is verified.
@@
Let $\theta\in[0,1]$ and let $\bar{x}\in\mathrm{Sol_{p}}(\theta)$. As $\bar{x}\in[0,1]$ and $G$ is therefore not active at $(\bar{x},\theta)$, we have $(\bar{x},0)\in\mathrm{Sol_{pd}}(\theta)$. It follows that
$D_{f}(\theta)$ $=\big\{u\in\mathbb{R}:\exists\,(x,\lambda)\in\mathrm{Sol_{pd}}(\theta),(0,u)\in\partial^{c}F(x,\theta)+\left(\lambda\nabla_{\!x}G(x,\theta),0\right)\big\}$
@@
Consider the absolutely continuous curve $\gamma\colon[0,1]\to\mathbb{R}$ defined as $t\mapsto t$. Now we suppose, for the sake of contradiction, that $f$ has the chain rule property for $D_{f}$. Then, for almost all $t\in[0,1]$, and for any $u\in D_{f}(\gamma(t))$, in particular any $u\in[-1,1]$,
@@
This is impossible. So $f$ does not have the chain rule property for $D_{f}$, i.e, $D_{f}$ is not conservative for $f$. ∎